OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y =____x +____ will fan and medal :)

OpenStudy (igreen):

First we plug it in point-slope form. \(y - y_1 = m(x - x_1)\) Where \(x_1\) is the x-value of the point, \(y_1\) is the y-value of the point, and \(m\) is the slope. So can you plug them in?

OpenStudy (anonymous):

y-4= 2(x-3) ?

OpenStudy (igreen):

Yep! \(y - 4 = 2(x - 3)\) Now we distribute 2 into the parenthesis: \(y - 4 = 2x - 6\) Now add 4 to both sides, what do you get?

OpenStudy (anonymous):

y=2x-10?

OpenStudy (igreen):

Actually check again..it would be: \(y = 2x - 6 + 4\)

OpenStudy (anonymous):

ohh

OpenStudy (igreen):

Now add -6 + 4

OpenStudy (anonymous):

-2

OpenStudy (igreen):

Yep, so our final answer is: y = 2x - 2

OpenStudy (anonymous):

but the answer has to be y= ___x + ___ @iGreen

OpenStudy (igreen):

It is.. y = 2x - 2 '2' goes in the first blank, and '-2' goes in the 2nd blank.

OpenStudy (anonymous):

so its basically 2x + -2

OpenStudy (igreen):

Yes.